I know the answer to this question but I would like someone to walk me through the steps of solving it/ what formula can be used to calculate systolic pressure:

WHAT IS THE SYSTOLIC PRESSURE OF A PERSON WITH A DIASTOLIC
PRESSURE OF 100 mm Hg AND A MAP OF 120 mm Hg
Answer #1

To calculate Systolic pressure we can modify the follwing formula,

MAP = SBP + 2(DP) / 3, so,

SBP = 3(MAP) - 2 (DP)

Given,

DP = 100 mm / hg,

MAP = 120 mm / hg,

SBP = 3 (120) - 2(100) = 160 mm / hg.